### v1

# PromptBench: A Unified Library for Evaluation of Large Language Models

Abstract
The evaluation of large language models (LLMs) is crucial to assess their performance and mitigate potential security risks. In this paper, we introduce PromptBench, a unified library to evaluate LLMs. It consists of several key components that are easily used and extended by researchers: prompt construction, prompt engineering, dataset and model loading, adversarial prompt attack, dynamic evaluation protocols, and analysis tools. PromptBench is designed to be an open, general, and flexible codebase for research purpose that can facilitate original study in creating new benchmarks, deploying downstream applications, and designing new evaluation protocols. The code is available at: https://github.com/microsoft/promptbench and will be continuously supported.
Keywords: Evaluation, large language models, framework
1 Introduction
Large language models (LLMs) have been revolutionizing numerous applications, impacting virtually every aspect of human life and society unprecedentedly. They are trained on a massive scale of carefully curated corpora (Brown et al., 2020; Team, 2023; Computer, 2023), enabling humans to interact with them easily and flexibly through natural language to solve complicated real-world tasks (OpenAI, 2023b) through prompts.
Evaluation is of paramount importance to understand the true capabilities of LLMs, mitigate potential risks, and eventually, benefit society further (Eisenstein, 2023; Chang et al., 2023). Recent efforts have evaluated LLMs from diverse aspects (Liang et al., 2022; Zheng et al., 2023; Li et al., 2023b; Huang et al., 2023). Among the findings, one of the most important is that current LLMs are sensitive to prompts (Wang et al., 2023b), vulnerable to adversarial prompt attacks (Zhu et al., 2023b), and exposed to data contamination (Willig et al., 2023; Zhou et al., 2023b; Zhu et al., 2023a), which pose severe security and privacy issues (Wang et al., 2023a; Simmons, 2022). Given the increasing popularity of LLMs, it is indispensable to develop a unified codebase to enable easy, fast, and flexible evaluation.
Existing libraries such as LlamaIndex (Liu, 2022), semantic kernel (Microsoft, 2023), and LangChain (Chase, 2022) offer an entire LLM pipeline for various downstream applications, they are not built inherently for evaluation purpose. LlamaIndex and LangChain enhance LLM applications by incorporating databases and various data sources, enabling advanced, context-aware functionalities. Semantic Kernel aims to merge AI services with programming languages for versatile AI app development. The eval-harness (Gao et al., 2023) library offers a comprehensive framework for evaluating generative language models, but it does not support other evaluations such as adversarial prompt attacks, prompt engineering, and dynamic evaluation. Consequently, there is an urgent need for a unified library dedicated to the comprehensive evaluation of diverse aspects, particularly for research purposes.
This paper introduces PromptBench, a unified python library for evaluating LLMs from comprehensive dimensions. [^1] The name “PromptBench” is the same as (Zhu et al., 2023b) where we only evaluated the robustness against adversarial prompt attack. We decided to keep the name and heavily extend that project. It consists of a wide range of LLMs and evaluation datasets, covering diverse tasks, evaluation protocols, adversarial prompt attacks, and prompt engineering techniques. As a holistic library, it also supports several analysis tools for interpreting the results. Our library is designed in a modular fashion, allowing researchers to easily build evaluation pipelines for their own projects. We open-source PromptBench with comprehensive documents and tutorials [^2] https://promptbench.readthedocs.io/en/latest/ to support easy, flexible, and collaborative evaluation. We believe PromptBench could boost our understanding of capabilities of LLMs and facilitate new research from the community.
2 PromptBench
PromptBench can be easily installed either via pip install promptbench or git clone. In this section, we briefly introduce the components of PromptBench and how to use it to build an evaluation pipeline for LLMs. An overview of PromptBench is shown in Figure 1.
Figure 1: The components and supported research areas of PromptBench.
2.1 Components
Models. PromptBench supports both open-source and proprietary LLMs and it is open to add more. Currently, it supports Flan-T5-large (Chung et al., 2022), Dolly (Databricks, 2023), Vicuna (Chiang et al., 2023), Llama2 (Touvron et al., 2023b), Cerebras-GPT (Dey et al., 2023), GPT-NEOX (Black et al., 2022), Flan-UL2 (Brain, 2023), phi-1.5 (Li et al., 2023c), PaLM2 (Anil et al., 2023), ChatGPT (OpenAI, 2023a), and GPT-4 (OpenAI, 2023b). PromptBench provides a unified LLMModel interface to allow easy construction and inference of a model with specified max generating tokens and generating temperature. More details of the supported models are shown in Appendix A.1.
Datasets and tasks. Currently, PromptBench consists of $1212$ diverse tasks with $2222$ public datasets and naturally supports more. It supports: sentiment analysis (SST-2 (Socher et al., 2013)), grammar correctness (CoLA (Warstadt et al., 2018)), duplicate sentence detection (QQP (Wang et al., 2017) and MRPC (Dolan and Brockett, 2005)), natural language inference (MNLI (Williams et al., 2018), QNLI (Wang et al., 2019), RTE (Wang et al., 2019), and WNLI (Levesque et al., 2012)), multi-task knowledge (MMLU dataset (Hendrycks et al., 2021)), reading comprehension (SQuAD V2 dataset (Rajpurkar et al., 2018)), translation (UN Multi(Eisele and Chen, 2010), IWSLT 2017 (Cettolo et al., 2017)), math problem-solving (Math (Saxton et al., 2019) and GSM8K(Cobbe et al., 2021)), logical reasoning (Boolean Expressions (bench authors, 2023)), commonsense reasoning (Commensense QA (Talmor et al., 2019), QASC (Khot et al., 2020), NummerSense (Lin et al., 2020), Date (bench authors, 2023) and Object Tracking (bench authors, 2023)), symbolic reasoning (LastLetterConcat (Wei et al., 2023)), algorithm (Valid Parentheses dataset (bench authors, 2023)) Through the unified DatasetLoader interface, it supports easy and customizable dataset loading and processing.
Prompts and prompt engineering. PromptBench offers a suite of $44$ distinct prompt types, and additionally, users have the flexibility to craft custom prompts using the Prompt interface. Task-oriented prompts are structured to clearly delineate the specific task expected of the model, whereas role-oriented prompts position the model in a defined role, such as an expert, advisor, or translator. These prompt categories are adaptable for both zero-shot and few-shot learning contexts, offering diverse application possibilities. Moreover, PromptBench currently includes $66$ prominent prompt engineering methods: few-shot Chain-of-Thought (Wei et al., 2023), zero-shot Chain-of-Thought (Kojima et al., 2022), EmotionPrompt (Li et al., 2023a), Expert Prompting (Xu et al., 2023), Generated Knowledge (Liu et al., 2022), and Least to Most (Zhou et al., 2023a). Our framework is not only equipped for the easy integration of these existing techniques through the prompt_engineering module but is also actively evolving to encompass a broader spectrum of prompt engineering methods, enhancing its adaptability in varied evaluation scenarios.
Adversarial prompt attacks. To facilitate the investigation of LLMs’ robustness on prompts, PromptBench integrates $77$ types of adversarial prompt attacks (Zhu et al., 2023b): TextBugger (Li et al., 2019), TextFooler (Jin et al., 2019), BertAttack (Li et al., 2020), DeepWordBug (Gao et al., 2018), Checklist (Ribeiro et al., 2020), StressTest (Naik et al., 2018), and semantics (Zhu et al., 2023b). These attacks can be easily called via the prompt_attack interface. It also supports the usage of curated adversarial prompts to efficiently evaluate the robustness of LLMs.
Different evaluation protocols. By default, PromptBench supports the standard protocol, i.e., the direct inference. PromptBench further supports dynamic (Zhu et al., 2023a) and semantic (Liu et al., 2023) evaluation protocols by dynamically generating testing data. It is open to integrate more new protocols to avoid data contamination.
Analysis tools. Finally, PromptBench offers a series of analysis tools to help researchers analyze their results. Particularly, it support sweep running to get the benchmark results. Then, attention visualization analysis can be done through the utils interface. PromptBench also supports word frequency analysis to analyze the words used in attacks as well as defense analysis by integrating word correction tools.
2.2 Evaluation pipeline
PromptBench allows easy construction of an evaluation pipeline via four steps. Firstly, specify task and then load dataset via pb.DatasetLoader. PromptBench offers a streamlined one-line API for loading the desired dataset. Secondly, users can customize LLMs using the pb.LLMModel, which provides integrated inference pipelines compatible with most LLMs implemented in Huggingface. Thirdly, the prompt for the specified dataset and task is defined via pb.Prompt. Users have the option to input a list of prompts for evaluation and performance comparison. In cases where no prompts are supplied, our default prompts for the dataset are utilized. Finally, the pipeline requires the definition of input and output processing functions via class InputProcess and class OutputProcess defined in pb.utils.dataprocess, as well as the evaluation function via pb.metrics. The detailed introduction of the components are shown in Appendix A.
2.3 Supported research topics
PromptBench is designed mainly for research purpose, thus it is easy to customize for different topics. As shown in Figure 1(b), it supports different evaluation topics from the research community including benchmarks, scenarios, and protocols. In benchmarks research, it supports standard natural language understanding, natural language generation, and reasoning tasks. It can also be extended to support research on AI agent and interdisciplinary study. In scenario research, it supports adversarial and out-of-distribution evaluation, and can also support other topics such as hallucination and bias by changing the metrics and DatasetLoader interface. In protocol, it naturally supports standard and dynamic evaluation, and can further be verified by including measurement theory.
PromptBench offers three leaderboards to allow easy comparison: adversarial prompt attack, prompt engineering, and dynamic evaluation, as shown in Appendix B. Researchers are welcome to submit new results to our platform. Extensibility is shown in Appendix C that allows convenient extension of the framework.
3 Conclusion
This work presented PromptBench, a unified framework for evaluation of large language models. The library is designed in modular fashion to allow users build evaluation pipeline by composing different models, tasks, and prompts. It also facilitates several research directions such as prompt engineering, adversarial prompt attacks, and dynamic evaluation. We treat PromptBench as the first step towards assessing the true capabilities and exploring the boundaries of current LLMs, and believe the benchmark and analysis results from our library could shed lights on designing more robust and human-aligned models. PromptBench is a long-term project and will be actively updated to fulfill the needs from the research community. We welcome contributions from all potential contributors.
Appendix A Details of PromptBench
A.1 Models
In this section, we list the LLMs implemented in PromptBench.
  • Flan-T5-large (Chung et al., 2022): Google’s Flan-T5-large, a variation of the Text-to-Text Transfer Transformer (T5).
  • Dolly-6B (Databricks, 2023): The Dolly-6B model, developed by Databricks, is a 6-billion parameter causal language model. It is an extension of EleutherAI’s GPT-J (Wang and Komatsuzaki, 2021), further refined with Stanford’s Alpaca (Taori et al., 2023) dataset comprising 52K question/answer pairs.
  • Vicuna series (Chiang et al., 2023): Developed from the LLaMA-13B base model, Vicuna-13B integrates over 70K user-shared conversations from ShareGPT.com, leveraging public APIs for data acquisition.
  • Cerebras series (Dey et al., 2023): Modeled on the GPT-3 architecture, Cerebras-13B is part of the Cerebras-GPT series, trained according to Chinchilla scaling laws (Hoffmann et al., 2022) to optimize computational efficiency.
  • Llama2 series (Touvron et al., 2023a): Engineered by Meta AI’s FAIR team, the Llama2 model is an autoregressive language model adopting the transformer architecture.
  • GPT-NEOX-20B (Black et al., 2022): This variant, part of the extensive GPT model series, features 20 billion parameters, exemplifying large-scale language model implementation.
  • Flan-UL2 (Brain, 2023): Flan-UL2, an encoder-decoder model, is grounded in the T5 architecture and enhanced with Flan prompt tuning and dataset techniques.
  • ChatGPT (OpenAI, 2023a) and GPT-4 (OpenAI, 2023b): OpenAI’s ChatGPT and GPT-4 are advanced iterations of the GPT series, with ChatGPT tailored for interactive tasks and GPT-4 being the most proficient in the series.
  • phi-1.5 (Li et al., 2023c): phi-1.5 is a transformer architecture with 1.3 billion parameters, builds upon the dataset used for phi-1 with the addition of diverse NLP synthetic texts.
  • PaLM 2 (Anil et al., 2023) PaLM 2 is an advanced language model that excels in multilingual and reasoning capabilities, offering greater computational efficiency than its predecessor, PaLM. This Transformer-based model enhances performance across various model sizes in English, multilingual tasks, and reasoning challenges.
A.2 Tasks and Datasets
  • GLUE (Wang et al., 2019) The GLUE benchmark (General Language Understanding Evaluation) offers a suite of tasks to evaluate the capability of NLP models in understanding language. For this research, we employed 8 specific tasks: Sentiment Analysis (SST-2 (Socher et al., 2013)), Grammar Correctness (CoLA (Warstadt et al., 2018)), Identifying Duplicate Sentences (QQP (Wang et al., 2017), MRPC (Dolan and Brockett, 2005)), and various Natural Language Inference tasks (MNLI (Williams et al., 2018), QNLI (Wang et al., 2019), RTE (Wang et al., 2019), WNLI (Levesque et al., 2012)).
  • MMLU (Hendrycks et al., 2021) The MMLU dataset tests the broad knowledge and problem-solving skills of large language models through 57 tasks with multiple-choice questions in fields like mathematics, history, and computer science. It is a comprehensive multitask benchmark.
  • SQuAD V2 (Rajpurkar et al., 2018) The SQuAD v2 dataset is pivotal in training and assessing NLP models for reading comprehension. It builds upon the original SQuAD by adding unanswerable questions, making it more challenging. Models must either identify the correct answer in the text or recognize questions as unanswerable.
  • UN Multi (Eisele and Chen, 2010) Comprising texts in the six official United Nations languages, the Multi UN dataset is a vast parallel corpus from UN documents. However, its focus on formal texts may restrict its use in informal or conversational language contexts.
  • IWSLT 2017 (Cettolo et al., 2017) Designed for spoken language translation system evaluation, the IWSLT 2017 dataset includes multilingual, multi-domain text data, primarily from the TED Talks Open Translation Project. It encompasses numerous language pairs, providing a rich resource for translation tasks.
  • Math (Saxton et al., 2019) The DeepMind Mathematics Dataset assesses AI models’ mathematical reasoning by posing a wide array of math problems, from algebra to calculus. It tests the models’ understanding and logical reasoning in mathematics.
  • BigBench (bench authors, 2023) The BIG-bench is a collaborative benchmark designed to evaluate the capabilities of large language models and predict their future potential. It consists of over 200 tasks, contributed by 444 authors from 132 institutions, covering a wide range of topics like linguistics, math, common-sense reasoning, and more. These tasks are intended to probe areas believed to be beyond the current capabilities of language models.
  • GSM8K (Cobbe et al., 2021) The GSM8K dataset is a collection of 8.5K high-quality, linguistically diverse grade school math word problems. It was created by human problem writers and is divided into 7.5K training problems and 1K test problems. These problems, which require 2 to 8 steps to solve, primarily involve basic arithmetic operations and are designed to be solvable by a bright middle school student.
  • CommonsenseQA (Talmor et al., 2019) The CommonsenseQA dataset is a challenging commonsense question-answering dataset. It comprises 12,247 questions with 5 multiple-choice answers each.
  • QASC (Khot et al., 2020) QASC (Question Answering via Sentence Composition) is a specialized collection designed for question-answering tasks with a focus on sentence composition. It comprises 9,980 eight-way multiple-choice questions about grade school science, divided into 8,134 for training, 926 for development, and 920 for testing .(In our evaluation, we use development part.) The dataset is notable for its emphasis on multi-hop reasoning, requiring the retrieval and composition of facts from a broad corpus to answer each question.
  • NummerSense (Lin et al., 2020) NumerSense is a unique numerical commonsense reasoning probing task, featuring a diagnostic dataset with 3,145 masked-word-prediction probes. This dataset has applications in tasks such as knowledge base completion and open-domain question answering.
A.3 Evaluation protocols
DyVal (Zhu et al., 2023a) is an approach for dynamic evaluation of LLMs by creating complexity-tailored evaluation samples on-the-fly, as opposed to relying on static benchmarks. DyVal synthesized seven distinct reasoning tasks, including: (1) Mathematics, focusing on arithmetic calculations and linear equation solving; (2) Logical Reasoning, involving boolean, deductive, and abductive logic; and (3) Algorithmic Analysis, covering reachability and the maximum sum path problem. MSTemp (Liu et al., 2023) stands for the semantic evalaution protocol which generate out-of-distribution samples by relying on evlauator LLMs and word replacement.
A.4 Prompts
A.4.1 Prompts
Our study examines four prompt categories, differentiated by their intended function and the required number of labeled samples. Task-oriented prompts are designed to clearly define the model’s task, prompting it to generate outputs relevant to the task using its inherent pre-training knowledge. In contrast, role-oriented prompts position the model as a particular entity, like an expert, advisor, or translator, thereby subtly guiding the expected output format and behavior through the assumed role. Both categories can be adapted for zero-shot and few-shot learning contexts. We randomly choose three training set examples for each task to form the few shot examples. Examples of various prompt types are illustrated in Table 1.
Table 1: Examples of $44$ types of prompts.
Zero shot Task oriented
Determine if the given pair of statements can be considered the same by responding with ’equivalent’ or ’not_equivalent’.
Role oriented
As an instrument for question comparison evaluation, consider the questions and determine if their meaning is the same, responding with ’equivalent’ for similar questions or ’not_equivalent’ for different questions.
Few shot Task oriented
Review the sentence below and identify whether its grammar is ’Acceptable’ or ’Unacceptable’: Here are three examples. Sentence: Our friends won’t buy this analysis, let alone the next one we propose. Answer: acceptable. Sentence: One more pseudo generalization and I’m giving up. Answer: acceptable. Sentence: They drank the pub. Answer: unacceptable.
Role oriented
Functioning as a grammar evaluation tool, analyze the given sentence and decide if it is grammatically correct, responding with ’acceptable’ or ’unacceptable’: Here are three examples. Sentence: Our friends won’t buy this analysis, let alone the next one we propose. Answer: acceptable. Sentence: One more pseudo generalization and I’m giving up. Answer: acceptable. Sentence: They drank the pub. Answer: unacceptable.
A.4.2 Prompt Engineering
Prompt engineering is a process of structuring and optimizing prompts to efficiently use AI models. Methods in prompt engineering , such as chain-of-thought(Wei et al., 2023), generated knowledge prompting(Liu et al., 2022) and so on, help improve reasoning ability and task performance of AI models. We implement 6 prominent prompt engineering methods:
  • Chain-of-Thought (Wei et al., 2023): This method involves breaking down complex, multi-step problems into smaller, intermediate steps, enabling Models to tackle more intricate reasoning tasks. Chain-of-Thought differs from standard few-shot prompting by not just providing questions and answers but prompting the model to produce intermediate reasoning steps before arriving at the final answer.
  • Zero-Shot Chain-of-Thought (Kojima et al., 2022): Zero-Shot Chain of Thought improves Chain of Thought by simplifying the prompting process. The key innovation in Zero-Shot Chain-of-Thought is appending the phrase “Let’s think step by step” to the end of a question.
  • EmotionPrompt (Li et al., 2023a): Drawing inspiration from psychology and social science theories about human emotional intelligence, this method adds emotional stimuli to origin prompts. For example: “This is very important to my career.”
  • Expert Prompting (Xu et al., 2023): The key idea is to let model be an expert in role playing. To generate the expert identity, we first provide several instruction-expert pair exemplars, then the model generates an expert identity of this question. Finally, we ask the model to answer the instruction conditioned on expert identity.
  • Generated Knowledge (Liu et al., 2022): Generated Knowledge Prompting is a method where a model first generates knowledge and then uses this generated information as additional input to answer questions. It enhances commonsense reasoning in AI without requiring task-specific knowledge integration or a structured knowledge base.
  • Least to Most (Zhou et al., 2023a): Least to Most breaks down a complex problem into a series of simpler subproblems and then solves them in sequence. The key idea is to solve each subproblem by using the answers to previously solved subproblems. This method is particularly useful for tasks that require solving problems harder than the exemplars shown in the prompts.
A.4.3 Adversarial Prompt Attacks
Adversarial prompt attacks, as proposed by Zhu et al. (2023b), aims to simulate potential disturbances that could naturally arise in practical scenarios. The proposed prompt attacks are intended to resemble common user errors or expressions, as users often make various mistakes when inputting prompts, such as typos, diverse word choices, and different sentence constructions. The prompt attacks encompass four distinct levels:
  • Character-level: Techniques such as TextBugger (Li et al., 2019) and DeepWordBug (Gao et al., 2018) are employed. These methods introduce errors or typos into words by altering characters.
  • Word-level: Attacks like BertAttack (Li et al., 2020) and TextFooler (Jin et al., 2019) are utilized. They focus on substituting words with their synonyms or contextually similar alternatives.
  • Sentence-level: StressTest (Naik et al., 2018) and CheckList (Ribeiro et al., 2020) are applied. These attacks add irrelevant or redundant sentences to prompts.
  • Semantic-level: To simulate the linguistic styles of different global regions.
A.5 Pipeline
The full pipeline of using PromptBench for evaluation is shown in Figure 2.
Figure 2: A pipeline for evaluation of LLMs.
Appendix B Benchmark Results
B.1 Adversarial prompt robustness
The partial results of the robustness of different models on a range of tasks are presented in Figure 3. All models exhibit vulnerability to adversarial prompts, with ChatGPT and GPT-4 demonstrating the strongest robustness.
Figure 3: Adversarial prompt robustness results.
Figure 4: Comparison among different prompt engineering techniques.
Figure 5: DyVal results.
B.2 Prompt engineering
Prompt engineering results are shown in Figure 4. Most methods are effective for special fields, so these methods can not surpass the baseline in every dataset.
B.3 Dynamic evaluation
Figure 5 illustrates the outcomes of dynamic evaluations across various models and tasks. GPT-4 outperforms its counterparts significantly, yet there remains potential for enhancement in the performance of linear equation, abductive logic, and max sum path task.
Appendix C Extensibility
Each module in PromptBench can be easily extended. In the following, we provide basic guidelines for customizing your own datasets, models, prompt engineering methods, and evaluation metrics.
C.1 Add new datasets
Adding new datasets involves two steps:
  1. 1.
    Implementing a New Dataset Class: Datasets are supposed to be implemented in dataload/dataset.py and inherit from the Dataset class. For your custom dataset, implement the __init__ method to load your dataset. We recommend organizing your data samples as dictionaries to facilitate the input process.
  2. 2.
    Adding an Interface: After customizing the dataset class, register it in the DataLoader class within dataload.py.
C.2 Add new models
Similar to adding new datasets, the addition of new models also consists of two steps.
  1. 1.
    Implementing a New Model Class: Models should be implemented in dataload/model.py, inheriting from the LLMModel class. In your customized model, you should implement self.tokenizer and self.model. You may also customize your own predict function for inference. If the predict function is not customized, the default predict function inherited from LLMModel will be used.
  2. 2.
    Adding an Interface: After customizing the model class, register it in the _create_model function within the class LLMModel in __init__.py.
C.3 Add new prompt engineering methods
Adding new methods in prompt engineering is similar to steps of C.1 and C.2.
  1. 1.
    Implementing a New Methods Class: Methods should be implemented in
    prompt_engineering Module. Firstly, create a new .py file for your methods. Then implement two functions: __init__ and query. For unified management, two points need be noticed: 1. all methods should inherits from Base class that has common code for prompt engineering methods. 2. prompts used in methods should be stored in prompts/method_oriented.py.
  2. 2.
    Adding an Interface: After implementing a new methods, register it in the METHOD_MAP that is used to map method names to their corresponding class.
C.4 Add new metrics and input/output process functions
New evaluation metrics should be implemented as static functions in class Eval within the metrics module. Similarly, new input/output process functions should be implemented as static functions in class InputProcess and class OutputProcess in the utils module.

### v2

@@
 In this paper, we introduce PromptBench, a unified library to evaluate LLMs.
-It consists of several key components that are easily used and extended by researchers: prompt construction, prompt engineering, dataset and model loading, adversarial prompt attack, dynamic evaluation protocols, and analysis tools.
-PromptBench is designed to be an open, general, and flexible codebase for research purpose that can facilitate original study in creating new benchmarks, deploying downstream applications, and designing new evaluation protocols.
+It consists of several key components that can be easily used and extended by researchers: prompt construction, prompt engineering, dataset and model loading, adversarial prompt attack, dynamic evaluation protocols, and analysis tools.
+PromptBench is designed as an open, general, and flexible codebase for research purpose.
+It aims to facilitate original study in creating new benchmarks, deploying downstream applications, and designing new evaluation protocols.
 The code is available at: https://github.com/microsoft/promptbench and will be continuously supported.
@@
 1 Introduction
-Large language models (LLMs) have been revolutionizing numerous applications, impacting virtually every aspect of human life and society unprecedentedly.
-They are trained on a massive scale of carefully curated corpora (Brown et al., 2020; Team, 2023; Computer, 2023), enabling humans to interact with them easily and flexibly through natural language to solve complicated real-world tasks (OpenAI, 2023b) through prompts.
+Large language models (LLMs) have been revolutionizing various aspects of human life and society unprecedentedly.
 Evaluation is of paramount importance to understand the true capabilities of LLMs, mitigate potential risks, and eventually, benefit society further (Eisenstein, 2023; Chang et al., 2023).
@@
 Among the findings, one of the most important is that current LLMs are sensitive to prompts (Wang et al., 2023b), vulnerable to adversarial prompt attacks (Zhu et al., 2023b), and exposed to data contamination (Willig et al., 2023; Zhou et al., 2023b; Zhu et al., 2023a), which pose severe security and privacy issues (Wang et al., 2023a; Simmons, 2022).
+On top of that, there have been various prompt learning algorithms developed based on different evaluation metrics, such as BDPL (Diao et al., 2022), GrIPS (Prasad et al., 2022) and Plum (Pan et al., 2023).
 Given the increasing popularity of LLMs, it is indispensable to develop a unified codebase to enable easy, fast, and flexible evaluation.
-Existing libraries such as LlamaIndex (Liu, 2022), semantic kernel (Microsoft, 2023), and LangChain (Chase, 2022) offer an entire LLM pipeline for various downstream applications, they are not built inherently for evaluation purpose.
-LlamaIndex and LangChain enhance LLM applications by incorporating databases and various data sources, enabling advanced, context-aware functionalities.
+There are existing libraries such as LlamaIndex (Liu, 2022), semantic kernel (Microsoft, 2023), and LangChain (Chase, 2022).
+However, LlamaIndex and LangChain enhance LLM applications by incorporating databases and various data sources, enabling context-aware functionalities.
 Semantic Kernel aims to merge AI services with programming languages for versatile AI app development.
-The eval-harness (Gao et al., 2023) library offers a comprehensive framework for evaluating generative language models, but it does not support other evaluations such as adversarial prompt attacks, prompt engineering, and dynamic evaluation.
+Eval-harness (Gao et al., 2023) offers a comprehensive framework for evaluating generative language models, but does not support other evaluations such as adversarial prompt attacks, prompt engineering, and dynamic evaluation.
+Zeno (Zeno, 2023) is an AI evaluation platform supporting interaction and visualization, but it is not easy to customize.
+LiteLLM (BerriAI, 2023) implements a unified API call for different LLM service prodiders, but it does not support research topics such as robustness, prompt engineering, and new protocols.
+Detailed comparisons are shown in Appendix A.
 Consequently, there is an urgent need for a unified library dedicated to the comprehensive evaluation of diverse aspects, particularly for research purposes.
@@
 We open-source PromptBench with comprehensive documents and tutorials [^2] https://promptbench.readthedocs.io/en/latest/ to support easy, flexible, and collaborative evaluation.
-We believe PromptBench could boost our understanding of capabilities of LLMs and facilitate new research from the community.
+We believe PromptBench could enhance our understanding of LLMs’ capabilities and spur new research within the community.
 2 PromptBench
@@
 PromptBench supports both open-source and proprietary LLMs and it is open to add more.
-Currently, it supports Flan-T5-large (Chung et al., 2022), Dolly (Databricks, 2023), Vicuna (Chiang et al., 2023), Llama2 (Touvron et al., 2023b), Cerebras-GPT (Dey et al., 2023), GPT-NEOX (Black et al., 2022), Flan-UL2 (Brain, 2023), phi-1.5 (Li et al., 2023c), PaLM2 (Anil et al., 2023), ChatGPT (OpenAI, 2023a), and GPT-4 (OpenAI, 2023b).
+Currently, it supports Flan-T5-large (Chung et al., 2022), Dolly (Databricks, 2023), Vicuna (Chiang et al., 2023), Llama2 series (Touvron et al., 2023b), Cerebras-GPT (Dey et al., 2023), GPT-NEOX (Black et al., 2022), Flan-UL2 (Brain, 2023), phi-1.5 (Li et al., 2023c), PaLM2 (Anil et al., 2023), ChatGPT (OpenAI, 2023a), and GPT-4 (OpenAI, 2023b).
 PromptBench provides a unified LLMModel interface to allow easy construction and inference of a model with specified max generating tokens and generating temperature.
-More details of the supported models are shown in Appendix A.1.
+More details of the supported models are shown in Appendix B.1.
 Datasets and tasks.
@@
 Finally, the pipeline requires the definition of input and output processing functions via class InputProcess and class OutputProcess defined in pb.utils.dataprocess, as well as the evaluation function via pb.metrics.
-The detailed introduction of the components are shown in Appendix A.
+The detailed introduction of the components are shown in Appendix B.
 2.3 Supported research topics
@@
 In protocol, it naturally supports standard and dynamic evaluation, and can further be verified by including measurement theory.
-PromptBench offers three leaderboards to allow easy comparison: adversarial prompt attack, prompt engineering, and dynamic evaluation, as shown in Appendix B.
+PromptBench offers three leaderboards to allow easy comparison: adversarial prompt attack, prompt engineering, and dynamic evaluation, as shown in Appendix C.
 Researchers are welcome to submit new results to our platform.
-Extensibility is shown in Appendix C that allows convenient extension of the framework.
+Extensibility is shown in Appendix D that allows convenient extension of the framework.
 3 Conclusion
-This work presented PromptBench, a unified framework for evaluation of large language models.
+We presented PromptBench, a unified framework for LLMs evaluation.
 The library is designed in modular fashion to allow users build evaluation pipeline by composing different models, tasks, and prompts.
@@
 We treat PromptBench as the first step towards assessing the true capabilities and exploring the boundaries of current LLMs, and believe the benchmark and analysis results from our library could shed lights on designing more robust and human-aligned models.
-PromptBench is a long-term project and will be actively updated to fulfill the needs from the research community.
+PromptBench is a long-term project and will be actively maintained.
 We welcome contributions from all potential contributors.
-Appendix A Details of PromptBench
-A.1 Models
+Appendix A Comparison with Related Code Libraries
+Table 1: Comparison with related code libraries.
+Library Purpose Customization Functions
+OpenAI Evals (Contributors, 2023b)
+A framework for evaluating LLMs or systems built using LLMs.
+model, dataset, prompt, eval
+Evaluation pipelines Benchmarks
+OpenCompass (Contributors, 2023c)
+One-stop platform for large model evaluation, aiming to provide a transparent benchmark for foundation models.
+model, dataset
+Evaluation pipelines Benchmarks Leaderboards
+promptfoo (Contributors, 2023a)
+A framework for evaluating prompts and large language models.
+model, dataset, prompt, eval
+Evaluation pipelines
+LM Evaluation Harness (Gao et al., 2023)
+A framework for evaluation of autoregressive LMs.
+model, dataset, prompt, eval
+Evaluation pipelines Benchmarks Leaderboards
+HELM (Liang et al., 2022)
+Holistic Evaluation of Language Models.
+model
+Evaluation pipelines Benchmark Leaderboard
+PromptBench (Ours)
+Research-focused evaluation toolkit.
+model, dataset, prompt, prompt engineering, eval
+Evaluation pipelines Prompt Engineering Prompt attacks Dynamic evaluation Leaderboard
+Appendix B Details of PromptBench
+B.1 Models
 In this section, we list the LLMs implemented in PromptBench.
+Open-source LLMs:
   • Flan-T5-large (Chung et al., 2022): Google’s Flan-T5-large, a variation of the Text-to-Text Transfer Transformer (T5).
@@
   • Flan-UL2 (Brain, 2023): Flan-UL2, an encoder-decoder model, is grounded in the T5 architecture and enhanced with Flan prompt tuning and dataset techniques.
+  • phi-1.5 and phi-2 (Li et al., 2023c): phi-1.5 is an LLM with 1.3 billion parameters, builds upon the dataset used for phi-1 with the addition of diverse NLP synthetic texts.
+  • Mistral 7B (Jiang et al., 2023): Mistral 7B is trained by Mistral AI team.
+It excels in tasks like reasoning, mathematics, and code generation.
+It uses grouped-query attention for faster inference and sliding window attention for efficient handling of sequences.
+There’s also an instruction-following version, Mistral 7B-Instruct.
+  • Mixtral8x7B (Mixtral, 2023): Engineering by Mistral AI team, this model is a high-quality sparse mixture of experts model (SMoE) with open weights.
+There’s also an instruction-following version, Mixtral 8x7B Instruct.
+  • Baichuan2 series (Yang et al., 2023): Baichuan2 is developed by Baichuan Intelligent.
+Trained on 2.6 trillion high-quality tokens, it achieves the best results in its size class on multiple authoritative benchmarks in Chinese, English, and multilingual general and domain-specific tasks.
+  • Yi series (Yi, 2023): Developed by 01.AI, the Yi series are next-generation open-source large language models.
+Trained on a 3T multilingual corpus, they excel in language understanding, commonsense reasoning, and reading comprehension.
+Proprietary LLMs:
   • ChatGPT (OpenAI, 2023a) and GPT-4 (OpenAI, 2023b): OpenAI’s ChatGPT and GPT-4 are advanced iterations of the GPT series, with ChatGPT tailored for interactive tasks and GPT-4 being the most proficient in the series.
-  • phi-1.5 (Li et al., 2023c): phi-1.5 is a transformer architecture with 1.3 billion parameters, builds upon the dataset used for phi-1 with the addition of diverse NLP synthetic texts.
-  • PaLM 2 (Anil et al., 2023) PaLM 2 is an advanced language model that excels in multilingual and reasoning capabilities, offering greater computational efficiency than its predecessor, PaLM.
+  • PaLM 2 (Anil et al., 2023): PaLM 2 is an advanced language model that excels in multilingual and reasoning capabilities, offering greater computational efficiency than its predecessor, PaLM.
 This Transformer-based model enhances performance across various model sizes in English, multilingual tasks, and reasoning challenges.
-A.2 Tasks and Datasets
-  • GLUE (Wang et al., 2019) The GLUE benchmark (General Language Understanding Evaluation) offers a suite of tasks to evaluate the capability of NLP models in understanding language.
+  • Gemini (Google, 2023): Gemini is Google’s new family of LLMs consisting of Gemini Nano, Gemini Pro, and Gemini Ultra. Gemini Pro is released via API service.
+B.2 Tasks and Datasets
+  • GLUE (Wang et al., 2019): The GLUE benchmark (General Language Understanding Evaluation) offers a suite of tasks to evaluate the capability of NLP models in understanding language.
 For this research, we employed 8 specific tasks: Sentiment Analysis (SST-2 (Socher et al., 2013)), Grammar Correctness (CoLA (Warstadt et al., 2018)), Identifying Duplicate Sentences (QQP (Wang et al., 2017), MRPC (Dolan and Brockett, 2005)), and various Natural Language Inference tasks (MNLI (Williams et al., 2018), QNLI (Wang et al., 2019), RTE (Wang et al., 2019), WNLI (Levesque et al., 2012)).
-  • MMLU (Hendrycks et al., 2021) The MMLU dataset tests the broad knowledge and problem-solving skills of large language models through 57 tasks with multiple-choice questions in fields like mathematics, history, and computer science.
+  • MMLU (Hendrycks et al., 2021): The MMLU dataset tests the broad knowledge and problem-solving skills of large language models through 57 tasks with multiple-choice questions in fields like mathematics, history, and computer science.
 It is a comprehensive multitask benchmark.
-  • SQuAD V2 (Rajpurkar et al., 2018) The SQuAD v2 dataset is pivotal in training and assessing NLP models for reading comprehension.
+  • SQuAD V2 (Rajpurkar et al., 2018): The SQuAD v2 dataset is pivotal in training and assessing NLP models for reading comprehension.
 It builds upon the original SQuAD by adding unanswerable questions, making it more challenging.
 Models must either identify the correct answer in the text or recognize questions as unanswerable.
-  • UN Multi (Eisele and Chen, 2010) Comprising texts in the six official United Nations languages, the Multi UN dataset is a vast parallel corpus from UN documents.
+  • UN Multi (Eisele and Chen, 2010): Comprising texts in the six official United Nations languages, the Multi UN dataset is a vast parallel corpus from UN documents.
 However, its focus on formal texts may restrict its use in informal or conversational language contexts.
-  • IWSLT 2017 (Cettolo et al., 2017) Designed for spoken language translation system evaluation, the IWSLT 2017 dataset includes multilingual, multi-domain text data, primarily from the TED Talks Open Translation Project.
+  • IWSLT 2017 (Cettolo et al., 2017): Designed for spoken language translation system evaluation, the IWSLT 2017 dataset includes multilingual, multi-domain text data, primarily from the TED Talks Open Translation Project.
 It encompasses numerous language pairs, providing a rich resource for translation tasks.
-  • Math (Saxton et al., 2019) The DeepMind Mathematics Dataset assesses AI models’ mathematical reasoning by posing a wide array of math problems, from algebra to calculus.
+  • Math (Saxton et al., 2019): The DeepMind Mathematics Dataset assesses AI models’ mathematical reasoning by posing a wide array of math problems, from algebra to calculus.
 It tests the models’ understanding and logical reasoning in mathematics.
-  • BigBench (bench authors, 2023) The BIG-bench is a collaborative benchmark designed to evaluate the capabilities of large language models and predict their future potential.
+  • BIG-Bench (bench authors, 2023): BIG-bench is a collaborative benchmark designed to evaluate the capabilities of large language models and predict their future potential.
 It consists of over 200 tasks, contributed by 444 authors from 132 institutions, covering a wide range of topics like linguistics, math, common-sense reasoning, and more.
-These tasks are intended to probe areas believed to be beyond the current capabilities of language models.
-  • GSM8K (Cobbe et al., 2021) The GSM8K dataset is a collection of 8.5K high-quality, linguistically diverse grade school math word problems.
+These tasks are intended to probe areas believed to be beyond the current capabilities of LMs.
+  • GSM8K (Cobbe et al., 2021): The GSM8K dataset is a collection of 8.5K high-quality, linguistically diverse grade school math word problems.
 It was created by human problem writers and is divided into 7.5K training problems and 1K test problems.
 These problems, which require 2 to 8 steps to solve, primarily involve basic arithmetic operations and are designed to be solvable by a bright middle school student.
-  • CommonsenseQA (Talmor et al., 2019) The CommonsenseQA dataset is a challenging commonsense question-answering dataset. It comprises 12,247 questions with 5 multiple-choice answers each.
-  • QASC (Khot et al., 2020) QASC (Question Answering via Sentence Composition) is a specialized collection designed for question-answering tasks with a focus on sentence composition.
+  • CommonsenseQA (Talmor et al., 2019): The CommonsenseQA dataset is a challenging commonsense question-answering dataset. It comprises 12,247 questions with 5 multiple-choice answers each.
+  • QASC (Khot et al., 2020): QASC (Question Answering via Sentence Composition) is a specialized collection designed for question-answering tasks with a focus on sentence composition.
 It comprises 9,980 eight-way multiple-choice questions about grade school science, divided into 8,134 for training, 926 for development, and 920 for testing .(In our evaluation, we use development part.) The dataset is notable for its emphasis on multi-hop reasoning, requiring the retrieval and composition of facts from a broad corpus to answer each question.
-  • NummerSense (Lin et al., 2020) NumerSense is a unique numerical commonsense reasoning probing task, featuring a diagnostic dataset with 3,145 masked-word-prediction probes.
+  • NummerSense (Lin et al., 2020): NumerSense is a unique numerical commonsense reasoning probing task, featuring a diagnostic dataset with 3,145 masked-word-prediction probes.
 This dataset has applications in tasks such as knowledge base completion and open-domain question answering.
-A.3 Evaluation protocols
+B.3 Evaluation protocols
 DyVal (Zhu et al., 2023a) is an approach for dynamic evaluation of LLMs by creating complexity-tailored evaluation samples on-the-fly, as opposed to relying on static benchmarks.
@@
 MSTemp (Liu et al., 2023) stands for the semantic evalaution protocol which generate out-of-distribution samples by relying on evlauator LLMs and word replacement.
-A.4 Prompts
-A.4.1 Prompts
+B.4 Prompts
+B.4.1 Prompts
 Our study examines four prompt categories, differentiated by their intended function and the required number of labeled samples.
@@
 Answer: unacceptable.
-A.4.2 Prompt Engineering
+B.4.2 Prompt Engineering
 Prompt engineering is a process of structuring and optimizing prompts to efficiently use AI models.
@@
 This method is particularly useful for tasks that require solving problems harder than the exemplars shown in the prompts.
-A.4.3 Adversarial Prompt Attacks
+Note that there are plenty of prompt engineering techniques and we tried our best to include those general techniques instead of specific prompt engineering techniques such as Tree of Thoughts (Yao et al., 2023) that requires specific prompt design and decomposition of each problem.
+B.4.3 Adversarial Prompt Attacks
 Adversarial prompt attacks, as proposed by Zhu et al.
@@
   • Semantic-level: To simulate the linguistic styles of different global regions.
-A.5 Pipeline
+B.5 Pipeline
 The full pipeline of using PromptBench for evaluation is shown in Figure 2.
 Figure 2: A pipeline for evaluation of LLMs.
-Appendix B Benchmark Results
-B.1 Adversarial prompt robustness
+Appendix C Benchmark Results
+C.1 Adversarial prompt robustness
 The partial results of the robustness of different models on a range of tasks are presented in Figure 3.
@@
 Figure 5: DyVal results.
-B.2 Prompt engineering
+C.2 Prompt engineering
 Prompt engineering results are shown in Figure 4. Most methods are effective for special fields, so these methods can not surpass the baseline in every dataset.
-B.3 Dynamic evaluation
+C.3 Dynamic evaluation
 Figure 5 illustrates the outcomes of dynamic evaluations across various models and tasks.
 GPT-4 outperforms its counterparts significantly, yet there remains potential for enhancement in the performance of linear equation, abductive logic, and max sum path task.
-Appendix C Extensibility
+Appendix D Extensibility
 Each module in PromptBench can be easily extended. In the following, we provide basic guidelines for customizing your own datasets, models, prompt engineering methods, and evaluation metrics.
-C.1 Add new datasets
+D.1 Add new datasets
 Adding new datasets involves two steps:
@@
     Adding an Interface: After customizing the dataset class, register it in the DataLoader class within dataload.py.
-C.2 Add new models
+D.2 Add new models
 Similar to adding new datasets, the addition of new models also consists of two steps.
   1. 1.
-    Implementing a New Model Class: Models should be implemented in dataload/model.py, inheriting from the LLMModel class.
+    Implementing a New Model Class: Models should be implemented in models/model.py, inheriting from the LLMModel class.
 In your customized model, you should implement self.tokenizer and self.model.
@@
   2. 2.
-    Adding an Interface: After customizing the model class, register it in the _create_model function within the class LLMModel in __init__.py.
-C.3 Add new prompt engineering methods
+    Adding an Interface: After customizing the model class, register it in the _create_model function within the class LLMModel and MODEL_LIST dictionary in __init__.py.
+D.3 Add new prompt engineering methods
 Adding new methods in prompt engineering is similar to steps of C.1 and C.2.
@@
     Adding an Interface: After implementing a new methods, register it in the METHOD_MAP that is used to map method names to their corresponding class.
-C.4 Add new metrics and input/output process functions
+D.4 Add new metrics and input/output process functions
 New evaluation metrics should be implemented as static functions in class Eval within the metrics module.

### v3

@@
 1 Introduction
-Large language models (LLMs) have been revolutionizing various aspects of human life and society unprecedentedly.
+Large language models (LLMs) are revolutionizing aspects of human life and society, such as medical diagnostics (McDuff et al., 2023; Thirunavukarasu et al., 2024), and educational tools (Ho et al., 2023).
 Evaluation is of paramount importance to understand the true capabilities of LLMs, mitigate potential risks, and eventually, benefit society further (Eisenstein, 2023; Chang et al., 2023).
-Recent efforts have evaluated LLMs from diverse aspects (Liang et al., 2022; Zheng et al., 2023; Li et al., 2023b; Huang et al., 2023).
-Among the findings, one of the most important is that current LLMs are sensitive to prompts (Wang et al., 2023b), vulnerable to adversarial prompt attacks (Zhu et al., 2023b), and exposed to data contamination (Willig et al., 2023; Zhou et al., 2023b; Zhu et al., 2023a), which pose severe security and privacy issues (Wang et al., 2023a; Simmons, 2022).
+Recent efforts have evaluated LLMs from diverse aspects (Hendrycks et al., 2021; Liang et al., 2022; Zheng et al., 2023; Li et al., 2023c; Huang et al., 2023; HuggingFace, 2023).
+Among the findings, one of the most important is that current LLMs are sensitive to prompts (Wang et al., 2023b), vulnerable to adversarial prompt attacks (Zhu et al., 2023b), and exposed to testset data contamination (Willig et al., 2023; Zhou et al., 2023b; Zhu et al., 2023a), which pose severe security and privacy issues (Wang et al., 2023a; Simmons, 2022).
 On top of that, there have been various prompt learning algorithms developed based on different evaluation metrics, such as BDPL (Diao et al., 2022), GrIPS (Prasad et al., 2022) and Plum (Pan et al., 2023).
-Given the increasing popularity of LLMs, it is indispensable to develop a unified codebase to enable easy, fast, and flexible evaluation.
+Given the increasing popularity of LLMs, it is indispensable to develop a unified codebase to enable easy, fast, and flexible evaluation of large foundation models.
 There are existing libraries such as LlamaIndex (Liu, 2022), semantic kernel (Microsoft, 2023), and LangChain (Chase, 2022).
-However, LlamaIndex and LangChain enhance LLM applications by incorporating databases and various data sources, enabling context-aware functionalities.
+LlamaIndex and LangChain enhance LLM applications by incorporating databases and various data sources.
 Semantic Kernel aims to merge AI services with programming languages for versatile AI app development.
-Eval-harness (Gao et al., 2023) offers a comprehensive framework for evaluating generative language models, but does not support other evaluations such as adversarial prompt attacks, prompt engineering, and dynamic evaluation.
+Eval-harness (Gao et al., 2023) offers a comprehensive framework for evaluating generative language models.
 Zeno (Zeno, 2023) is an AI evaluation platform supporting interaction and visualization, but it is not easy to customize.
-LiteLLM (BerriAI, 2023) implements a unified API call for different LLM service prodiders, but it does not support research topics such as robustness, prompt engineering, and new protocols.
-Detailed comparisons are shown in Appendix A.
-Consequently, there is an urgent need for a unified library dedicated to the comprehensive evaluation of diverse aspects, particularly for research purposes.
-This paper introduces PromptBench, a unified python library for evaluating LLMs from comprehensive dimensions.
-[^1] The name “PromptBench” is the same as (Zhu et al., 2023b) where we only evaluated the robustness against adversarial prompt attack.
-We decided to keep the name and heavily extend that project.
+LiteLLM (BerriAI, 2023) implements a unified API call for different LLM service prodiders.
+This paper introduces PromptBench, a unified Python library to evaluate LLMs from comprehensive dimensions.
+It is designed to fill the gaps current libraries have, offering comprehensive support not only for standard model evaluations but also for advanced scenarios including adversarial prompt attacks and dynamic evaluations.
+Its extensible architecture allows for the incorporation of new evaluation protocols, addressing the limitations found in other tools.
+The detailed comparisons are shown in Appendix A, highlighting how PromptBench provides a more complete toolkit for the nuanced evaluation of language models, especially in research contexts where adaptability and thoroughness are paramount.
 It consists of a wide range of LLMs and evaluation datasets, covering diverse tasks, evaluation protocols, adversarial prompt attacks, and prompt engineering techniques.
@@
 We open-source PromptBench with comprehensive documents and tutorials [^1] https://promptbench.readthedocs.io/en/latest/ to support easy, flexible, and collaborative evaluation.
-We believe PromptBench could enhance our understanding of LLMs’ capabilities and spur new research within the community.
+We believe PromptBench could enhance our understanding of LLMs and spur new research within the community.
 2 PromptBench
@@
 Models.
-PromptBench supports both open-source and proprietary LLMs and it is open to add more.
-Currently, it supports Flan-T5-large (Chung et al., 2022), Dolly (Databricks, 2023), Vicuna (Chiang et al., 2023), Llama2 series (Touvron et al., 2023b), Cerebras-GPT (Dey et al., 2023), GPT-NEOX (Black et al., 2022), Flan-UL2 (Brain, 2023), phi-1.5 (Li et al., 2023c), PaLM2 (Anil et al., 2023), ChatGPT (OpenAI, 2023a), and GPT-4 (OpenAI, 2023b).
-PromptBench provides a unified LLMModel interface to allow easy construction and inference of a model with specified max generating tokens and generating temperature.
+PromptBench supports both open-source and proprietary LLMs and VLMs and it is open to add more.
+Currently, it supports a diverse range of LLMs and VLMs, ranging from Llama2 series (Touvron et al., 2023b), Mixtral series (Jiang et al., 2024), LlaVa series (Liu et al., 2023a) to GPT series (OpenAI, 2023a, b).
+PromptBench provides unified LLMModel and VLMModel interfaces to allow easy construction and inference of a model with specified max generating tokens and generating temperature.
+The interfaces also support customized models, including those that have been fine-tuned for specific applications.
 More details of the supported models are shown in Appendix B.1.
 Datasets and tasks.
-Currently, PromptBench consists of $1212$ diverse tasks with $2222$ public datasets and naturally supports more.
-It supports: sentiment analysis (SST-2 (Socher et al., 2013)), grammar correctness (CoLA (Warstadt et al., 2018)), duplicate sentence detection (QQP (Wang et al., 2017) and MRPC (Dolan and Brockett, 2005)), natural language inference (MNLI (Williams et al., 2018), QNLI (Wang et al., 2019), RTE (Wang et al., 2019), and WNLI (Levesque et al., 2012)), multi-task knowledge (MMLU dataset (Hendrycks et al., 2021)), reading comprehension (SQuAD V2 dataset (Rajpurkar et al., 2018)), translation (UN Multi(Eisele and Chen, 2010), IWSLT 2017 (Cettolo et al., 2017)), math problem-solving (Math (Saxton et al., 2019) and GSM8K(Cobbe et al., 2021)), logical reasoning (Boolean Expressions (bench authors, 2023)), commonsense reasoning (Commensense QA (Talmor et al., 2019), QASC (Khot et al., 2020), NummerSense (Lin et al., 2020), Date (bench authors, 2023) and Object Tracking (bench authors, 2023)), symbolic reasoning (LastLetterConcat (Wei et al., 2023)), algorithm (Valid Parentheses dataset (bench authors, 2023)) Through the unified DatasetLoader interface, it supports easy and customizable dataset loading and processing.
+PromptBench comprises a wide array of tasks, currently supporting diverse challenges across $1212$ tasks and $2222$ public datasets, with the capacity for additional expansions.
+The supported tasks include fundamental NLP tasks such as sentiment analysis, grammar correctness, and duplicate sentence detection, as well as complex challenges involving natural language inference, multi-task knowledge, and reading comprehension.
+It also covers specialized areas like translation, mathematical problem-solving, and various forms of reasoning—logical, commonsense, symbolic, and algorithmic.
+For detailed descriptions of each dataset and specific task configurations (see Appendix B.2).
+The unified DatasetLoader interface facilitates easy and customizable loading and processing of these datasets, enhancing the usability and flexibility of PromptBench.
 Prompts and prompt engineering.
@@
 These prompt categories are adaptable for both zero-shot and few-shot learning contexts, offering diverse application possibilities.
-Moreover, PromptBench currently includes $66$ prominent prompt engineering methods: few-shot Chain-of-Thought (Wei et al., 2023), zero-shot Chain-of-Thought (Kojima et al., 2022), EmotionPrompt (Li et al., 2023a), Expert Prompting (Xu et al., 2023), Generated Knowledge (Liu et al., 2022), and Least to Most (Zhou et al., 2023a).
+Moreover, PromptBench currently includes $66$ prominent prompt engineering methods, details can be found in Appendix B.4.2.
 Our framework is not only equipped for the easy integration of these existing techniques through the prompt_engineering module but is also actively evolving to encompass a broader spectrum of prompt engineering methods, enhancing its adaptability in varied evaluation scenarios.
 Adversarial prompt attacks.
-To facilitate the investigation of LLMs’ robustness on prompts, PromptBench integrates $77$ types of adversarial prompt attacks (Zhu et al., 2023b): TextBugger (Li et al., 2019), TextFooler (Jin et al., 2019), BertAttack (Li et al., 2020), DeepWordBug (Gao et al., 2018), Checklist (Ribeiro et al., 2020), StressTest (Naik et al., 2018), and semantics (Zhu et al., 2023b).
-These attacks can be easily called via the prompt_attack interface.
-It also supports the usage of curated adversarial prompts to efficiently evaluate the robustness of LLMs.
+To facilitate the investigation of LLMs’ robustness on prompts, PromptBench integrates $44$ types of attacks: (1) character-level attacks (Li et al., 2019; Gao et al., 2018), which manipulate texts by introducing typos or errors to words; (2) word-level attacks (Jin et al., 2019; Li et al., 2020), which aim to replace words with synonyms or contextually similar words to deceive LLMs; (3) sentence-level attacks (Ribeiro et al., 2020; Naik et al., 2018), which append irrelevant or extraneous sentences to the end of prompts, intending to distract LLMs; (4) semantic-level attacks (Zhu et al., 2023b), which simulate the linguistic behavior of people from different countries.
+Details can be found in Appendix B.4.3 These attacks can be easily called via the prompt_attack interface.
+It also supports the usage of curated adversarial prompts to efficiently evaluate robustness.
 Different evaluation protocols.
 By default, PromptBench supports the standard protocol, i.e., the direct inference.
-PromptBench further supports dynamic (Zhu et al., 2023a) and semantic (Liu et al., 2023) evaluation protocols by dynamically generating testing data.
+PromptBench further supports dynamic (Zhu et al., 2023a) and semantic (Liu et al., 2023b) evaluation protocols by dynamically generating testing data.
 It is open to integrate more new protocols to avoid data contamination.
@@
 2.3 Supported research topics
-PromptBench is designed mainly for research purpose, thus it is easy to customize for different topics.
+Compared to current evaluation libraries, PromptBench is designed mainly for research purpose, thus it is easy to customize for different topics.
 As shown in Figure 1(b), it supports different evaluation topics from the research community including benchmarks, scenarios, and protocols.
@@
 Extensibility is shown in Appendix D that allows convenient extension of the framework.
-3 Conclusion
-We presented PromptBench, a unified framework for LLMs evaluation.
-The library is designed in modular fashion to allow users build evaluation pipeline by composing different models, tasks, and prompts.
-It also facilitates several research directions such as prompt engineering, adversarial prompt attacks, and dynamic evaluation.
-We treat PromptBench as the first step towards assessing the true capabilities and exploring the boundaries of current LLMs, and believe the benchmark and analysis results from our library could shed lights on designing more robust and human-aligned models.
-PromptBench is a long-term project and will be actively maintained.
-We welcome contributions from all potential contributors.
+3 Conclusion and Discussion
+We presented PromptBench, a unified framework for LLMs evaluation, designed in a modular fashion to build evaluation pipelines with various models, tasks, and prompts.
+It supports research in prompt engineering, adversarial attacks, and dynamic evaluation.
+PromptBench is the first step in assessing and exploring the capabilities of current LLMs.
+We believe our benchmark and analysis will inform the design of more robust, human-aligned models.
+In the future, new datasets, evaluation protocols, prompt variations, and analytical tools will be added into PromptBench.
+We also welcome any contributions for PromptBench.
+Despite its versatility, PromptBench has limitations.
+It may not cover all evaluation scenarios, and some metrics might miss nuanced performance differences.
+The framework’s effectiveness depends on the quality and diversity of datasets and prompts.
+Addressing these limitations is a priority for our ongoing and future work.
 Appendix A Comparison with Related Code Libraries
@@
 Library Purpose Customization Functions
-OpenAI Evals (Contributors, 2023b)
-A framework for evaluating LLMs or systems built using LLMs.
-model, dataset, prompt, eval
-Evaluation pipelines Benchmarks
-OpenCompass (Contributors, 2023c)
-One-stop platform for large model evaluation, aiming to provide a transparent benchmark for foundation models.
-model, dataset
-Evaluation pipelines Benchmarks Leaderboards
-promptfoo (Contributors, 2023a)
-A framework for evaluating prompts and large language models.
-model, dataset, prompt, eval
-Evaluation pipelines
-LM Evaluation Harness (Gao et al., 2023)
-A framework for evaluation of autoregressive LMs.
-model, dataset, prompt, eval
-Evaluation pipelines Benchmarks Leaderboards
-HELM (Liang et al., 2022)
-Holistic Evaluation of Language Models.
-model
-Evaluation pipelines Benchmark Leaderboard
-PromptBench (Ours)
-Research-focused evaluation toolkit.
-model, dataset, prompt, prompt engineering, eval
-Evaluation pipelines Prompt Engineering Prompt attacks Dynamic evaluation Leaderboard
+OpenAI Evals (Contributors, 2023b) A framework for evaluating LLMs or systems built using LLMs. model, dataset, prompt, eval Evaluation pipelines Benchmarks
+OpenCompass (Contributors, 2023c) One-stop platform for large model evaluation, aiming to provide a transparent benchmark for foundation models.
+model, dataset Evaluation pipelines Benchmarks Leaderboards
+promptfoo (Contributors, 2023a) A framework for evaluating prompts and large language models. model, dataset, prompt, eval Evaluation pipelines
+LM Evaluation Harness (Gao et al., 2023) A framework for evaluation of autoregressive LMs. model, dataset, prompt, eval Evaluation pipelines Benchmarks Leaderboards
+HELM (Liang et al., 2022) Holistic Evaluation of Language Models. model Evaluation pipelines Benchmark Leaderboard
+PromptBench (Ours) Research-focused evaluation toolkit. model, dataset, prompt, prompt engineering, eval Evaluation pipelines Prompt Engineering Prompt attacks Dynamic evaluation Leaderboard
 Appendix B Details of PromptBench
 B.1 Models
-In this section, we list the LLMs implemented in PromptBench.
+In this section, we list the LLMs and VLMS implemented in PromptBench.
 Open-source LLMs:
@@
   • Flan-UL2 (Brain, 2023): Flan-UL2, an encoder-decoder model, is grounded in the T5 architecture and enhanced with Flan prompt tuning and dataset techniques.
-  • phi-1.5 and phi-2 (Li et al., 2023c): phi-1.5 is an LLM with 1.3 billion parameters, builds upon the dataset used for phi-1 with the addition of diverse NLP synthetic texts.
+  • phi-1.5 and phi-2 (Li et al., 2023d): phi-1.5 is an LLM with 1.3 billion parameters, builds upon the dataset used for phi-1 with the addition of diverse NLP synthetic texts.
   • Mistral 7B (Jiang et al., 2023): Mistral 7B is trained by Mistral AI team.
@@
 Trained on a 3T multilingual corpus, they excel in language understanding, commonsense reasoning, and reading comprehension.
+  • BLIP2 (Li et al., 2023b): This visual-language model is proposed by Junnan Li, Dongxu Li, Silvio Savarese, Steven Hoi.
+BLIP-2 leverages frozen pre-trained image encoders and large language models (LLMs) by training a lightweight, 12-layer Transformer encoder in between them, achieving excellent performance in various vision-language tasks
+  • LLaVA (Liu et al., 2024): LlaVA (Language-Image LLaMA) is a multimodal model combining language and image data.
+It extends the LLaMA architecture to handle both modalities, enabling tasks like image captioning, visual question answering, and image-based text generation.
+  • Qwen-VL series(Bai et al., 2023): Qwen-VL (Qwen Large Vision Language Model) is the multimodal version of the large model series, Qwen (abbr.
+Tongyi Qianwen), proposed by Alibaba Cloud.
+Qwen-VL accepts image, text, and bounding box as inputs, outputs text, and bounding box.
+  • InternLM-XComposer2-VL (Dong et al., 2024): InternLM-XComposer2 is a cutting-edge vision-language model excelling in free-form text-image composition and comprehension, crafting content from diverse inputs like outlines, detailed specs, and reference images.
+Using a Partial LoRA (PLoRA) approach, it balances vision understanding and text composition.
 Proprietary LLMs:
-  • ChatGPT (OpenAI, 2023a) and GPT-4 (OpenAI, 2023b): OpenAI’s ChatGPT and GPT-4 are advanced iterations of the GPT series, with ChatGPT tailored for interactive tasks and GPT-4 being the most proficient in the series.
+  • ChatGPT (OpenAI, 2023a) and GPT-4 (OpenAI, 2023b): OpenAI’s ChatGPT and GPT-4 are advanced iterations of the GPT series.
+ChatGPT is tailored for interactive tasks, while GPT-4 is the most proficient in the series and supports image input.
   • PaLM 2 (Anil et al., 2023): PaLM 2 is an advanced language model that excels in multilingual and reasoning capabilities, offering greater computational efficiency than its predecessor, PaLM.
 This Transformer-based model enhances performance across various model sizes in English, multilingual tasks, and reasoning challenges.
-  • Gemini (Google, 2023): Gemini is Google’s new family of LLMs consisting of Gemini Nano, Gemini Pro, and Gemini Ultra. Gemini Pro is released via API service.
+  • Gemini (Google, 2023): The Gemini model is a multimodal language model developed by Google AI, capable of extracting insights from a diverse array of data formats, including images, and video.
 B.2 Tasks and Datasets
@@
 This dataset has applications in tasks such as knowledge base completion and open-domain question answering.
+  • VQAv2 (Goyal et al., 2017): Visual Question Answering (VQA) v2.0 is a dataset containing open-ended questions about images.
+These questions require an understanding of vision, language and commonsense knowledge to answer.
+It is the second version of the VQA dataset.
+  • NoCaps (Agrawal et al., 2019): NoCaps is a benchmark dataset for image captioning models that can describe images containing novel objects from object detection datasets.
+It consists of 166,100 human-generated captions describing 15,100 images from the Open Images validation and test sets.
+  • MMMU (Yue et al., 2023): MMMU is a comprehensive benchmark designed to evaluate multimodal models on massive multi-discipline tasks demanding college-level subject knowledge and deliberate reasoning.
+MMMU includes 11.5K meticulously collected multimodal questions from college exams, quizzes, and textbooks.
+These questions span 30 subjects and 183 subfields, comprising 32 highly heterogeneous image types, such as charts, diagrams, maps, tables, music sheets, and chemical structures.
+  • MathVista (Lu et al., 2023): MathVista is a comprehensive benchmark for mathematical reasoning in visual contexts.
+It includes three new datasets: IQTest, FunctionQA, and PaperQA.
+These datasets cover various visual domains and are designed to evaluate logical reasoning on puzzle test figures, algebraic reasoning using functional plots, and scientific reasoning with academic paper figures, respectively.
+  • AI2D (Kembhavi et al., 2016): AI2D is a dataset of illustrative diagrams for research on diagram understanding and associated question answering.
+It contains 5000 grade-school science diagrams with over 150,000 rich annotations, their ground truth syntactic parses, and more than 15,000 corresponding multiple-choice questions.
+  • ChartQA (Masry et al., 2022): ChartQA is a large-scale dataset of complex reasoning questions over charts that involve visual and logical operations, covering 9.6K human-written questions as well as 23.1K questions generated from human-written chart summaries.
+  • ScienceQA (Lu et al., 2022): ScienceQA is a large-scale dataset for multimodal reasoning with diverse science topics and annotations of answers, lectures and explanations.
+It covers 26 topics, 127 categories and 379 skills from natural science, language science, and social science and so on.
 B.3 Evaluation protocols
@@
 DyVal synthesized seven distinct reasoning tasks, including: (1) Mathematics, focusing on arithmetic calculations and linear equation solving; (2) Logical Reasoning, involving boolean, deductive, and abductive logic; and (3) Algorithmic Analysis, covering reachability and the maximum sum path problem.
-MSTemp (Liu et al., 2023) stands for the semantic evalaution protocol which generate out-of-distribution samples by relying on evlauator LLMs and word replacement.
+MSTemp (Liu et al., 2023b) stands for the semantic evalaution protocol which generate out-of-distribution samples by relying on evlauator LLMs and word replacement.
 B.4 Prompts
@@
 Table 2: Examples of $44$ types of prompts.
-Zero shot Task oriented
-Determine if the given pair of statements can be considered the same by responding with ’equivalent’ or ’not_equivalent’.
-Role oriented
-As an instrument for question comparison evaluation, consider the questions and determine if their meaning is the same, responding with ’equivalent’ for similar questions or ’not_equivalent’ for different questions.
-Few shot Task oriented
-Review the sentence below and identify whether its grammar is ’Acceptable’ or ’Unacceptable’: Here are three examples.
+Zero shot Task oriented Determine if the given pair of statements can be considered the same by responding with ’equivalent’ or ’not_equivalent’.
+Role oriented As an instrument for question comparison evaluation, consider the questions and determine if their meaning is the same, responding with ’equivalent’ for similar questions or ’not_equivalent’ for different questions.
+Few shot Task oriented Review the sentence below and identify whether its grammar is ’Acceptable’ or ’Unacceptable’: Here are three examples.
 Sentence: Our friends won’t buy this analysis, let alone the next one we propose.
@@
 Answer: unacceptable.
-Role oriented
-Functioning as a grammar evaluation tool, analyze the given sentence and decide if it is grammatically correct, responding with ’acceptable’ or ’unacceptable’: Here are three examples.
+Role oriented Functioning as a grammar evaluation tool, analyze the given sentence and decide if it is grammatically correct, responding with ’acceptable’ or ’unacceptable’: Here are three examples.
 Sentence: Our friends won’t buy this analysis, let alone the next one we propose.
@@
 Appendix D Extensibility
-Each module in PromptBench can be easily extended. In the following, we provide basic guidelines for customizing your own datasets, models, prompt engineering methods, and evaluation metrics.
+Each module in PromptBench can be easily extended.
+In the following, we provide basic guidelines for customizing your own datasets, models, prompt engineering methods, and evaluation metrics.
+The sample code for adding new datasets, models can be found in https://github.com/microsoft/promptbench/blob/main/examples/add_new_modules.md
 D.1 Add new datasets